OpenStudy (anonymous):

Divide (x^3-27)/(x-3)

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^3 - b^3 = (a - b)(a^2 + ab + b^2)}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^3 - 3^3 = (x - 3)(x^2 + 3x + 9)}\)

Parth (parthkohli):

Do you get it? @annej

Parth (parthkohli):

And then, something cancels out. You may even factor the result after the cancelling is done.

OpenStudy (anonymous):

ok